The perimeter of a rectangle is 34 feet, the diagonal is 13 feet. What is the length and the width?
Answer by edjones(8007)

The diagonal divides it into 2 equal right triangles.
a^2+b^2=c^2
b=17-a for one triangle
a^2+(17-a)^2=13^2
a^2+a^2-34a+289=169
2a^2-34a+120=0
a^2-17a+60=0
(a-12)(a-5)=0
a=12', b=5' length and width.
.
Ed